OpenStudy (anonymous):

6 less than the product of 13 and y

OpenStudy (anonymous):

6-13*y

OpenStudy (anonymous):

damn uXD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thx for the help guys

geerky42 (geerky42):

Shouldn't it be other way around? 13*y-6

OpenStudy (anonymous):

it doesnt matter cuz of pemdas

geerky42 (geerky42):

It does matter... \(13y-6\) is not same as \(6-13y\)

OpenStudy (anonymous):

yeah it should be 13*y-6

OpenStudy (anonymous):

yeah your right, thx for catching the mistake

OpenStudy (anonymous):

yes cuz you do the multiplication first then the subtraction because of Parentheses, exponents, multiplication, division, addition, and subtraction

OpenStudy (anonymous):

itll be the same either way

geerky42 (geerky42):

No, see; \(6-4 = 2\) \(4-6 = -2\) \(2\neq-2\)

geerky42 (geerky42):

In subtraction, order matters.

geerky42 (geerky42):

Same for division.

OpenStudy (anonymous):

we dont know what y is howd did you get 4?

geerky42 (geerky42):

It's just a example...

OpenStudy (anonymous):

ok then 13x4 is 52, 52-6 is 46

OpenStudy (anonymous):

13(3)-6=33 , 6-39= -33 difference

geerky42 (geerky42):

Ok, if it is other way around; \(6-13\times4\)... \(6-52=-46\)

OpenStudy (anonymous):

oh i see what your saying my fault i was think it was the same as addition

OpenStudy (anonymous):

my fault

geerky42 (geerky42):

Ok cool.